OpenStudy (anonymous):

help? :) Solve the following system using substitution. x+3y=8 2x-y=2

OpenStudy (australopithecus):

x =8 - 3y sub this into the other equation and solve for y

OpenStudy (australopithecus):

when I say other equation I mean 2x - y = 2 once you find y just sub the y value back into one of the equations and solve for x

OpenStudy (australopithecus):

any questions ask

OpenStudy (anonymous):

thankyou! im just really confused on how you even got to that point? ^^^

OpenStudy (australopithecus):

I just rearranged the equation to equal to one of the variables in this case x so x+3y=8 = -3y + x + 3y = 8 - 3y = x = 8 - 3y now that we know what x is equal to we can sub it into the other equation and solve for y

OpenStudy (australopithecus):

understand?

OpenStudy (australopithecus):

after subbing in you get 2(8-3y) - y = 2 notice now that there is only one variable in the equation and therefore you can solve for it. If you know what value y equals to you can use it to find x by plugging it into one of the original equations

OpenStudy (australopithecus):

Doing this we are finding the point where the two functions (lines) intersect